John walks 10 miles at average rate of 2 miles per hour and returns on a bicycle at an average rate of 10 miles per hour.how long (to the nearest hour)does the entire trip takes him?
Answer by rothauserc(4718)

use the formula
:
Rate * Time = Distance, then
:
Time = Distance/Rate
:
T1 = 10/2 = 5 hours
:
T2 = 10/10 = 1 hour
:
The entire trip takes 5 + 1 = 6 hours
: